The force on a wire carrying 7.75A is a maximum of1.83N when placed between the pole faces of a magnet.

Part A

If the pole faces are 66.5cm in diameter, what is the approximate strength of the magnetic field?

Express your answer to three significant figures and include the appropriate units.
Answer #1

F = I * L * B Sin theta

1.83 = 7.75 * 0.665 * B

B = 1.83 / (7.75 * 0.665)

B = 0.355 T